a student is trying to solve the equation (x + 10)(x - 10)=7. The student has set x + 7 and x - 10 = 7 and found that two solutions x=3, x=17. The student checks the results by plugging inthe solutions into the original equation and finds they do not work. How would you advise the student?
Answer by checkley77(12844)

(x + 10)(x - 10)=7 EXPAND THE LEFT HAND TERMS.
X^2-100=7 COMBINE LIKE TERMS.
X^2-100+7=0
X^2-97=0
X^2=97
X=SQRT97
X=+-9.85 ANS.